# Higher-order Sobolev and Rellich inequalities via iterated Talenti’s principle

### v1

@@ 4. Rellich inequalities and improvements
The proof follows a similar line of thoughts as the proof of Theorem 1.1: We observe that it is enough to prove it for smooth positive functions $u$ on bounded domains $\Omega$. In addition, we define $f=(-\Delta_{g})^{k}u$ and consider the problem (14), and its symmetrization (15), whose solution is denoted by $v$.
@@
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{T}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\prod_{s=1}^{j}\ell_{[s]}^{2}\left(\frac{|x|}{R}\right)\,\mathrm{d}v_{g}\right).$
  • Ghoussoub and Moradifam [12, Corollary 3.8/1]: Let $n\geq 5$, $\Omega\subseteq M$ be bounded, $k\geq 1$ be an integer, $R$ denote the radius of $\Omega^{\star}$ and define $r=R\cdot\exp_{[k-1]}(e)$. Then for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{k}\int_{\Omega}\frac{u^{2}}{|x|^{4}}\left(\prod_{i=1}^{j}\log_{[i]}\left(\frac{r}{|x|}\right)\right)^{-2}\,\mathrm{d}v_{g}\right).$
  • Gazzola, Grunau, and Mitidieri [10, Theorem 2]: Let $p\geq 2$, $n\geq 2p$, and $\Omega\subset M$ be bounded, then for every $u\in C_{0}^{\infty}(\Omega)$ one has
@@
  • Gazzola, Grunau, and Mitidieri [10, Corollary 2]: Let $\Omega\subset M$ be bounded and $m\geq 1$ be an integer with $n>2m$. Then there exist some constants $c_{1},c_{2},\dots,c_{m}$ depending only on $m$, $n$, and $\operatorname{vol}_{g}(\Omega)$ such that for every $u\in C_{0}^{\infty}(\Omega)$ one has

### v2

@@ 4. Rellich inequalities and improvements
The proof follows a similar line of thoughts as the proof of Theorem 1.1: We observe that it is enough to prove it for regular domains $\Omega$. In addition, we define $f=(-\Delta_{g})^{k}u$ and consider the problem (14), and its symmetrization (15), whose solution is denoted by $v$.
@@
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{\rho^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{T}\int_{\Omega}\frac{u^{2}}{\rho^{4}}\prod_{s=1}^{j}\ell_{[s]}^{2}\left(\frac{\rho}{R}\right)\,\mathrm{d}v_{g}\right).$
  • Ghoussoub and Moradifam [18, Corollary 3.8/1]: Let $n\geq 5$, $\Omega\subseteq M$ be regular, $k\geq 1$ be an integer, $R$ denote the radius of $\Omega^{\star}$ and define $r=R\cdot\exp_{[k-1]}(e)$. Then for every $u\in C_{0}^{\infty}(\Omega)$ one has
    $\int_{\Omega}|\Delta_{g}u|^{2}\,\mathrm{d}v_{g}\geq(c_{g})^{4}\cdot\left(\frac{n^{2}(n-4)^{2}}{16}\int_{\Omega}\frac{u^{2}}{\rho^{4}}\,\mathrm{d}v_{g}+\left(1+\frac{n(n-4)}{8}\right)\sum_{j=1}^{k}\int_{\Omega}\frac{u^{2}}{\rho^{4}}\left(\prod_{i=1}^{j}\log_{[i]}\left(\frac{r}{\rho}\right)\right)^{-2}\,\mathrm{d}v_{g}\right).$
  • Gazzola, Grunau, and Mitidieri [16, Theorem 2]: Let $p\geq 2$, $n\geq 2p$, and $\Omega\subset M$ be regular, then for every $u\in C_{0}^{\infty}(\Omega)$ one has
@@
  • Gazzola, Grunau, and Mitidieri [16, Corollary 2]: Let $\Omega\subset M$ be regular and $m\geq 1$ be an integer with $n>2m$. Then there exist some constants $c_{1},c_{2},\dots,c_{m}$ depending only on $m$, $n$, and $\operatorname{vol}_{g}(\Omega)$ such that for every $u\in C_{0}^{\infty}(\Omega)$ one has
@@
Acknowledgement. The authors would like to thank to the anonymous reviewers for carefully reading the article and providing valuable comments.